# Speeding up quantum measurement using space-time trade-off

### v2

@@ I Introduction
II Case study: Poisson measurement statistics
@@
In an idealised situation, the number of photons counted will follow a Poisson distribution
$\mathcal{P}_{\ket{j},t}^{(1)}(k)=L_{\nu_{j}(t)}(k),$ (1)
where $L_{\omega}(k)=\frac{\omega^{k}}{k!}e^{-\omega}$. Recall the mean and variance of a Poisson distribution coincide and are both equal, $\langle K\rangle_{L_{\omega}}=\Delta K^{2}_{L_{\omega}}=\omega$. The two distributions, $\mathcal{P}_{\ket{0},t}^{(1)}$ and $\mathcal{P}_{\ket{1},t}^{(1)}$, will have an overlap which is the source of error in measurement quality when assigning outcomes; no threshold value can be defined which will give you a perfect measurement by correctly discriminating the two states.
If $k_{1}$ photons are emitted during time $t_{1}$ and $k_{2}$ photons are emitted during a subsequent time $t_{2}$ such that the total number of photons emitted in time $t=t_{1}+t_{2}$ is $k=k_{1}+k_{2}$, then by the fundamental convolution property of the Poisson distribution,
$L_{\nu_{j}(t_{1}+t_{2})}(k)$ $=\sum_{k=k_{1}+k_{2}}L_{\nu_{j}(t_{1})}(k_{1})L_{\nu_{j}(t_{2})}(k_{2}),$
$=(L_{\nu_{j}(t_{1})}*L_{\nu_{j}(t_{2})})(k),$
$=L_{\nu_{j}(t_{1})+\nu_{j}(t_{2})}(k),$
i.e. the Poisson distribution convolves to itself. This implies that $\nu_{j}(t)=\mu_{j}t$, for a rate $\mu_{j}$ independent of $t$. In particular the convolution property means that $\nu_{j}(2t)=2\nu_{j}(t)$.
@@
III Exact space-time trade-off in the idealised Poisson case
@@
For now, for clarity of presentation, we analyse the case of independent detectors for each qubit, but come back to this point below. We denote by $k_{1}$ and $k_{2}$ the number of photons detected from the first and second qubit detectors respectively and we input the state |$\ket{\psi}=\ket{j}$ into our scheme. The measurement statistics for the pair of outcomes $(k_{1},k_{2})$, given we are measuring the post-CNOT state |$\ket{jj}$ for a time $t$, are
|$\mathcal{R}_{\ket{jj},t}(k_{1},k_{2})=L_{\mu_{j}t}(k_{1})L_{\mu_{j}t}(k_{2}).$ (4)
The outcomes $(k_{1},k_{2})$ form a pair of independent random variables, hence the probability distribution of their sum $k=k_{1}+k_{2}$ is the convolution of the individual probability distributions, thus ${\mathcal{P}_{\ket{j},t}^{(2)}(k)=(L_{\mu_{j}t}*L_{\mu_{j}t})(k)=L_{2\mu_{j}t}(k)}$.
An important feature is present here; since the probability distribution only considers the total sum of emissions from both fluorescing qubits, we can consider the distribution as describing the number of photons counted from a single detector which is collecting all the emitted photons, rather than collecting separately and then adding.
@@
IV General space-time trade-off
@@
We will assume that we have a measurement procedure that is time dependent, meaning that the duration of the measurement can be changed, which will change the quality of the measurement. Let us assume that the statistics of this measurement (with outcome $x$) when measuring the two basis states for a time $t$ is given by $\mathcal{P}_{\ket{j},t}^{(1)}(x)$, where the superscript reminds us that this is without using the measurement scheme, i.e. we are performing a single measurement. For this measurement we can calculate the SNR (2) between these two distributions. Let us assume the dependence with time, to leading order, is
$S(\mathcal{P}_{\ket{0},t}^{(1)},\mathcal{P}_{\ket{1},t}^{(1)})\sim t^{\alpha}.$ (7)
For the case of Poisson statistics above we saw that $\alpha~{}=~{}1/2$ however we may imagine an alternative dependence on time.
@@
Figure 2: Space-time trade-off for idealised and realistic devices. The parameters used for the plots are experimentally relevant and are taken from [28]. In plots a) and b), we use Poisson distributed measurement statistics (5) with $\mu_{0}=3.5\text{ms}^{-1}$ and $\mu_{1}=14\text{ms}^{-1}$ for different numbers of qubits $N$. a) SNR (2) against time from $N=1$ to $N=5$ qubits. The SNR always increases with time and qubit number and continues the trend beyond what is shown on the graph. b) time of measurement $t$ against number of qubits from $N=1$ to $N=10$ qubits for the SNR values on the grey dashed lines in a). Inset: same points but for the inverse time $1/t$, demonstrating the linear relationship (6). c) and d) show similar results but for realistic devices. In this case, a higher maximum distinguishability can be achieved by introducing more qubits and faster measurements can be made while also improving distinguishability. These plots are generated using the noise and decoherence model (IX), found in the Appendix. We use input parameters $\lambda=0.0041\text{ms}^{-1}$ and $p=0.01$ where $p$ is chosen as a conservative estimate of CNOT failure rate. In c), the SNR increases with time up to a maximum value then decreases due to decoherence. The rate at which the SNR improves increases with qubit number, as does the peak SNR achieved. Inset: points in the inset no longer follow a straight line, indicating that the space-time trade-off is not linear with qubit number $N$.
@@
In the above we saw that, in an idealised version of the scheme involving perfect CNOTs, by bringing in $N-1$ additional ancilliary qubits, applying CNOTs and measuring each, that the statistics of the sum of outcomes is precisely the convolution of the statistics of an individual measurement, as given in (5). Here, while general distributions $\mathcal{P}_{\ket{j},t}^{(1)}$ do not convolve to distributions of the same form, for perfect CNOTs we nevertheless have the fundamental property that the expectation and (due to independence) variance add,
$\langle X\rangle_{\mathcal{P}_{\ket{j},t}^{(N)}}$ $=N\langle X\rangle_{\mathcal{P}_{\ket{j},t}^{(1)}},$ $\Delta X^{2}_{\mathcal{P}_{\ket{j},t}^{(N)}}$ $=N\Delta X^{2}_{\mathcal{P}_{\ket{j},t}^{(1)}}.$
It therefore follows that the SNR, using the scheme, will be
@@
Thus, as long as $t^{\alpha}\sqrt{N}$ is constant, then the quality of the measurement is constant, and hence we still obtain a space-time trade-off. In particular, if $t_{S}^{(1)}$ is the duration of the measurement which achieves an SNR of $S$, then using the scheme, we can speed this up to a measurement time of $t^{(N)}_{S}$ given by
$t_{S}^{(1)}/t^{(N)}_{S}=N^{1/(2\alpha)}.$ (9)
This recovers (6) in the case $\alpha=1/2$ but shows more generally, if $\alpha>1/2$, we will still achieve a trade-off between space and time using this scheme, albeit not as good as (6). Nevertheless, if $\alpha<1/2$, we will achieve a “super-improvement” where the trade-off is better than (6).
V Space-time trade-off in more realistic settings
For our scheme to be useful in practical applications it is crucial to show that the speed-up is robust to realistic imperfection. To this end, we now describe a simple noise model to take into account the dominant sources of error. Since ancilla qubits are initialised in state |$\ket{0}$, imperfections in the CNOT gate operations can be modelled by considering errors on state |$\ket{10}$ being mapped to state |$\ket{00}$ with error probability p (see eg. [29]). Additionally, we consider imperfections in qubit readout caused by qubits in state |$\ket{1}$ falling into |$\ket{0}$ during the measurement, such as due to depuming by off-resonant scattering in atomic fluorescence measurements [30]. This error is included by applying a probabilistic decay at constant rate $\lambda$ during the measurement. See the Appendix for further details on our noise model.
The effects are summarised in Figs. 2c and 2d. The SNR is calculated using (26) in the Supplementary Material, which gives the SNR of a realistic $N$ qubit scheme entirely in terms of the expectation and variance of the number of photons counted for an $N=1$ qubit scheme $K_{\ket{j},t}^{(1)}$ and the expectation and variance of the CNOT noise probability distribution. We see that we still obtain a significant space-time trade-off, allowing for significantly faster measurements, even in the presence of noise, demonstrating the robustness of this scheme.
We note that our discussions have been framed in terms of the SNR, but we find essentially the same space-time trade-off if we consider the measurement infidelity: this is discussed in the Supplementary Material. In the perfect case this is shown analytically while for the imperfect case we present numerical results analogous to Fig. 2c and 2d.
@@ VI Conclusion
Our case study was framed around state dependent fluorescence qubit measurements which are particularly appropriate to trapped-ion and cold-atom qubits. But, we would like to reiterate that our ideas are much more general. For example, our ideas also apply, when using multiple detectors, to superconducting qubits. We give a very brief discussion of this in the Supplementary Material. We also believe, just as in the earlier discussions, that we can embody our scheme for superconductors using a single detector; this is the subject of future work.
@@
VII Note added
During preparation of this manuscript we became aware of the independent work of Petrosyan, Norrell, Poole and Saffman [31] which contains related ideas.
@@ VIII Acknowledgements
IX Appendix on noise model
We consider a noise model with two key ingredients: (i) Noise in the CNOT gates applied; (ii) decoherence of the individual systems. We outline here these two elements separately.
@@ IX Appendix on noise model
$\widetilde{C_{p}}(\ket{10}\bra{10})$ $=p\ket{00}\bra{00}+(1-p)\ket{11}\bra{11}.$ | . (10)
Note that since we are interested in the SNR as our figure of merit, we only need to consider the input states |$\ket{\psi}=\ket{0}$ and |$\ket{\psi}=\ket{1}$ of the system qubit. Furthermore, we assume our ancillary qubits are initialised perfectly in |$\ket{0}$ so we are only interested in how the noisy CNOT gates will act on the states |$\ket{00}$ and |$\ket{10}$. Thus (IX) defines the action on all states that occur in the scheme.
@@
Figure 3: Increase in peak SNR from Fig. 2c and 2d. This figure is extra detail for Figs. 2c and 2d and uses the same parameters. We plot the peak SNR $S_{max}$ for $N=1$ to $N=10$ qubits. While more qubits results in a higher peak SNR, this trend flattens out for higher qubit numbers before eventually dropping. The slight unevenness in the plot is an odd/even effect of the cascade CNOT compilation, see Supplementary Material.
@@
There are numerous ways to compile the gates and different compilations give different probability distributions for the number of qubits in state |$\ket{1}$, given an input state |$\ket{\psi}=\ket{1}$. These compilations are explored in more detail in the Supplementary Material. Let us denote by $\mathcal{T}^{(N)}_{p}(q)$ the probability distribution of the random variable $Q^{(N)}_{p}$ defining the number of qubits in state |$\ket{1}$ after applying the noisy CNOT gates. Any qubits not in state |$\ket{1}$ are in state |$\ket{0}$. Combining this with the probability distribution for $N$ qubits emitting $k$ photons (5), we find the resultant distribution using our scheme with noisy CNOT gates to be
@@
Our complete noise and decoherence model consists of combining the noisy CNOT protocol (IX) with the decaying Poisson distribution (12). Altogether, we find that the probability distributions for the number of photons emitted, using an $N$ qubit scheme, is given by
@@
$\mathcal{P}_{\ket{1},t}^{(N)}(k)$ $=\sum^{N}_{q=0}\mathcal{T}^{(N)}_{p}(q)(\mathcal{W}_{\mu_{1},\mu_{0},\lambda,t}^{*q}*L_{(N-q)\mu_{0}t})(k).$ (13)
These are used to produce the results in Figs. 2c and 2d where we use the cascade CNOT compilation, found in the Supplementary Material, for $\mathcal{T}^{(N)}_{p}(q)$.

### v3

# main_supplementary_material

@@ Abstract
.2 Case study: Poisson measurement statistics
@@
In an idealised situation, the number of photons counted will follow a Poisson distribution ${\mathcal{P}_{\ket{j},t}^{(1)}(k)=L_{\nu_{j}(t)}(k)}$, where $L_{\omega}(k)=\frac{\omega^{k}}{k!}e^{-\omega}$. Recall the mean and variance of a Poisson distribution coincide and are both equal, $\langle K\rangle_{L_{\omega}}=\Delta K^{2}_{L_{\omega}}=\omega$. The two distributions, $\mathcal{P}_{\ket{0},t}^{(1)}$ and $\mathcal{P}_{\ket{1},t}^{(1)}$, will have an overlap which is the source of error in measurement quality when assigning outcomes; no threshold value can be defined which will give you a perfect measurement by correctly discriminating the two states.
If $k_{1}$ photons are emitted during time $t_{1}$ and $k_{2}$ photons are emitted during a subsequent time $t_{2}$ such that the total number of photons emitted in time $t=t_{1}+t_{2}$ is $k=k_{1}+k_{2}$, then by the fundamental convolution property of the Poisson distribution, ${L_{\nu_{j}(t_{1}+t_{2})}(k)=(L_{\nu_{j}(t_{1})}*L_{\nu_{j}(t_{2})})(k)=L_{\nu_{j}(t_{1})+\nu_{j}(t_{2})}(k)}$, i.e. the Poisson distribution convolves to itself. This implies that $\nu_{j}(t)=\mu_{j}t$, for a rate $\mu_{j}$ independent of $t$. In particular the convolution property means that $\nu_{j}(2t)=2\nu_{j}(t)$.
@@
.3 Exact space-time trade-off in the idealised Poisson case
@@
For now, for clarity of presentation, we analyse the case of independent detectors for each qubit, but come back to this point below. We denote by $k_{1}$ and $k_{2}$ the number of photons detected from the first and second qubit detectors respectively and we input the state |$\ket{\psi}=\ket{j}$ into our scheme. The measurement statistics for the pair of outcomes $(k_{1},k_{2})$, given we are measuring the post-CNOT state |$\ket{jj}$ for a time $t$, are |${\mathcal{R}_{\ket{jj},t}(k_{1},k_{2})=L_{\mu_{j}t}(k_{1})L_{\mu_{j}t}(k_{2})}$. The outcomes $(k_{1},k_{2})$ form a pair of independent random variables, hence the probability distribution of their sum $k=k_{1}+k_{2}$ is the convolution of the individual probability distributions, thus ${\mathcal{P}_{\ket{j},t}^{(2)}(k)=(L_{\mu_{j}t}*L_{\mu_{j}t})(k)=L_{2\mu_{j}t}(k)}$. An important feature is present here; since the probability distribution only considers the total sum of emissions from both fluorescing qubits, we can consider the distribution as describing the number of photons counted from a single detector which is collecting all the emitted photons, rather than collecting separately and then adding.
@@
.4 General space-time trade-off
@@
We will assume that we have a measurement procedure that is time dependent, meaning that the duration of the measurement can be changed, which will change the quality of the measurement. Let us assume that the statistics of this measurement (with outcome $x$) when measuring the two basis states for a time $t$ is given by $\mathcal{P}_{\ket{j},t}^{(1)}(x)$, where the superscript reminds us that this is without using the measurement scheme, i.e. we are performing a single measurement. For this measurement we can calculate the SNR (1) between these two distributions. Let us assume the dependence with time, to leading order, is ${S(\mathcal{P}_{\ket{0},t}^{(1)},\mathcal{P}_{\ket{1},t}^{(1)})\sim t^{\alpha}.}$ For the case of Poisson statistics above we saw that $\alpha~{}=~{}1/2$ however we may imagine an alternative dependence on time.
@@
Figure 2: Space-time trade-off for idealised and realistic devices. The parameters used for the plots are experimentally relevant and are taken from [28]. In plots a) and b), we use Poisson distributed measurement statistics (3) with $\mu_{0}=3.5\text{ms}^{-1}$ and $\mu_{1}=14\text{ms}^{-1}$ for different numbers of qubits $N$. a) SNR (1) against time from $N=1$ to $N=5$ qubits. The SNR always increases with time and qubit number and continues the trend beyond what is shown on the graph. b) time of measurement $t$ against number of qubits from $N=1$ to $N=10$ qubits for the SNR values on the grey dashed lines in a). Inset: same points but for the inverse time $1/t$, demonstrating the linear relationship (4). c) and d) show similar results but for realistic devices. In this case, a higher maximum distinguishability can be achieved by introducing more qubits and faster measurements can be made while also improving distinguishability. These plots are generated using the noise and decoherence model (I.1), found in the Appendix. We use input parameters $\lambda=0.0041\text{ms}^{-1}$ and $p=0.01$ where $p$ is chosen as a conservative estimate of CNOT failure rate. In c), the SNR increases with time up to a maximum value then decreases due to decoherence. The rate at which the SNR improves increases with qubit number, as does the peak SNR achieved. Inset: points in the inset no longer follow a straight line, indicating that the space-time trade-off is not linear with qubit number $N$.
@@
In the above we saw that, in an idealised version of the scheme involving perfect CNOTs, by bringing in $N-1$ additional ancilliary qubits, applying CNOTs and measuring each, that the statistics of the sum of outcomes is precisely the convolution of the statistics of an individual measurement, as given in (3). Here, while general distributions $\mathcal{P}_{\ket{j},t}^{(1)}$ do not convolve to distributions of the same form, for perfect CNOTs we nevertheless have the fundamental property that the expectation and (due to independence) variance add, ${\langle X\rangle_{\mathcal{P}_{\ket{j},t}^{(N)}}=N\langle X\rangle_{\mathcal{P}_{\ket{j},t}^{(1)}}}$, ${\Delta X^{2}_{\mathcal{P}_{\ket{j},t}^{(N)}}=N\Delta X^{2}_{\mathcal{P}_{\ket{j},t}^{(1)}}}$. It therefore follows that the SNR, using the scheme, will be
@@
Thus, as long as $t^{\alpha}\sqrt{N}$ is constant, then the quality of the measurement is constant, and hence we still obtain a space-time trade-off. In particular, if $t_{S}^{(1)}$ is the duration of the measurement which achieves an SNR of $S$, then using the scheme, we can speed this up to a measurement time of $t^{(N)}_{S}$ given by ${t_{S}^{(1)}/t^{(N)}_{S}=N^{1/(2\alpha)}}$. This recovers (4) in the case $\alpha=1/2$ but shows more generally, if $\alpha>1/2$, we will still achieve a trade-off between space and time using this scheme, albeit not as good as (4). Nevertheless, if $\alpha<1/2$, we will achieve a “super-improvement” where the trade-off is better than (4).
.5 Space-time trade-off in more realistic settings
For our scheme to be useful in practical applications it is crucial to show that the speed-up is robust to realistic imperfection. To this end, we now describe a simple noise model to take into account the dominant sources of error. Since ancilla qubits are initialised in state |$\ket{0}$, imperfections in the CNOT gate operations can be modelled by considering errors on state |$\ket{10}$ being mapped to state |$\ket{00}$ with error probability $p$ (see eg. [29]). Additionally, we consider imperfections in qubit readout caused by qubits in state |$\ket{1}$ falling into |$\ket{0}$ during the measurement, such as due to depumping by off-resonant scattering in atomic fluorescence measurements [30]. This error is included by applying a probabilistic decay at constant rate $\lambda$ during the measurement. See the Appendix for further details on our noise model.
The effects are summarised in Figs. 2c and 2d. We see that we still obtain a significant space-time trade-off, allowing for significantly faster measurements, even in the presence of noise, demonstrating the robustness of this scheme.
Fig. 2 was generated using (I.2), in the Appendix, which gives a general expression for the SNR after using our scheme. It is applicable to essentially any arbitrary time-dependent noisy measurement process and accounts for arbitrary CNOT imperfections. We used (I.2) with the noise model outlined immediately above but, by using (I.2), the analogue of Fig. 2 could be reproduced much more generally for any specific architecture.
We note that our discussions have been framed in terms of the SNR, but we find essentially the same space-time trade-off if we consider the measurement infidelity: this is discussed in the Supplemental Material. In the perfect case this is shown analytically while for the imperfect case we present numerical results analogous to Fig. 2c and 2d.
@@
Our case study was framed around state dependent fluorescence qubit measurements which are particularly appropriate to trapped-ion and cold-atom qubits. We would like to reiterate that our ideas are much more general, as demonstrated by equation (I.2) in the Appendix. For example, our ideas also apply, when using multiple detectors, to superconducting qubits. We give a very brief discussion of this in the Supplemental Material. We also believe, just as in the earlier discussions, that we can embody our scheme for superconductors using a single detector; this is the subject of future work.
We have presented our results under the assumption that the CNOT duration is negligibly short compared to the measurement duration. The extent to which this is true is hardware specific, but it is straightforward to incorporate CNOT duration and, if it is sufficiently short, see that a trade-off will still be obtained. The relatively short two-qubit gate times compared to readout times in the following references demonstrate the advantage of our scheme to superconducting [31, 32] and neutral atom [1, 33] quantum computers while for trapped-ion quantum computers [34, 35] the advantage is less clear cut at present. We note however that this comparison is only relevant at the time of writing. Due to the fast development of quantum technologies the precise relation between two-qubit gate and readout durations is highly likely to change in the foreseeable future. Our view is that hardware providers will need to use every tool available to make their machines as effective as possible and our scheme has the potential to become a standard part of this toolbox.
One of the most exciting applications of our scheme is to mid-circuit measurements, such as those used in error correction. As an important example, some form of parity measurement is key to many error correction schemes: here the parity of the qubits that is needed can be found by coupling those qubits to an ancilla using CNOTs, then measuring the ancilla. This ancilla measurement, and hence the whole error detection, can be sped-up by our scheme.
@@
Note added.—-During preparation of this manuscript we became aware of the independent work of Petrosyan, Norrell, Poole and Saffman [36] which contains related ideas.
@@ I Acknowledgements
I.1 Appendix A: On noise model
We consider a noise model with two key ingredients: (i) Noise in the CNOT gates applied; (ii) decoherence of the individual systems. Initially we consider simple, albeit well-motivated models for both, but later in the Appendix we explain how the analysis can be extended to very general noise models. We first outline here these two ingredients separately in the simple model case.
@@ I.1 Appendix A: On noise model
$\widetilde{C_{p}}(\ket{10}\bra{10})$ $=p\ket{00}\bra{00}+(1-p)\ket{11}\bra{11}.$ | . (A1)
Note that since we are interested in the SNR as our figure of merit, we only need to consider the input states |$\ket{\psi}=\ket{0}$ and |$\ket{\psi}=\ket{1}$ of the system qubit. Furthermore, we assume our ancillary qubits are initialised perfectly in |$\ket{0}$ so we are only interested in how the noisy CNOT gates will act on the states |$\ket{00}$ and |$\ket{10}$. Thus (I.1) defines the action on all states that occur in the scheme.
@@
Figure 3: Increase in peak SNR from Fig. 2c and 2d. This figure is extra detail for Figs. 2c and 2d and uses the same parameters. We plot the peak SNR $S_{max}$ for $N=1$ to $N=10$ qubits. While more qubits results in a higher peak SNR, this trend flattens out for higher qubit numbers before eventually dropping. The slight unevenness in the plot is an odd/even effect of the cascade CNOT compilation, see the Supplemental Material.
@@
There are numerous ways to compile the gates and different compilations give different probability distributions for the number of qubits in state |$\ket{1}$, given an input state |$\ket{\psi}=\ket{1}$. These compilations are explored in more detail in the Supplemental Material. Let us denote by $\mathcal{T}^{(N)}_{p}(q)$ the probability distribution of the random variable $Q^{(N)}_{p}$ defining the number of qubits in state |$\ket{1}$ after applying the noisy CNOT gates. Any qubits not in state |$\ket{1}$ are in state |$\ket{0}$. Combining this with the probability distribution for $N$ qubits emitting $k$ photons (3), we find the resultant distribution using our scheme with noisy CNOT gates to be
@@
Our complete noise and decoherence model consists of combining the noisy CNOT protocol (I.1) with the decaying Poisson distribution (A3). Altogether, we find that the probability distributions for the number of photons emitted, using an $N$ qubit scheme, is given by
@@
$\mathcal{P}_{\ket{1},t}^{(N)}(k)$ $=\sum^{N}_{q=0}\mathcal{T}^{(N)}_{p}(q)(\mathcal{W}_{\mu_{1},\mu_{0},\lambda,t}^{*q}*L_{(N-q)\mu_{0}t})(k).$ (A4)
These are used to produce the results in Figs. 2c and 2d where we use the cascade CNOT compilation, found in the Supplemental Material, for $\mathcal{T}^{(N)}_{p}(q)$.
I.2 Appendix B: On general space-time trade-off
We now show how the scheme can be extended to very general noise/decoherence models with minimal assumptions. We allow for the CNOT gates to induce an arbitrary trace preserving quantum channel (i.e. we assume that the system does not leave the space spanned by the states |$\ket{0}$ and |$\ket{1}$ for each qubit) and we also assume that the measurement statistics are identical for each qubit. The SNR then depends only on (i) $P_{\ket{j},t}(k)$, the probability distribution of the random variable $K_{\ket{j},t}$ defining the measurement statistics of the basis state |$\ket{j}$ at time $t$ for an arbitrary time-dependent noisy measurement process and (ii) $T_{\ket{j}}^{(N)}(q)$, the probability distribution of the random variable $Q_{\ket{j}}^{(N)}$ defining the final number of qubits in state |$\ket{j}$ after applying the noisy CNOT gates to the input state |$\ket{j}$ and $N-1$ ancillary qubits initialised in state |$\ket{0}$. We note that the SNR will only depend on $T_{\ket{j}}^{(N)}(q)$, which is a limited statistic of the entangling CNOT process, so in order to compute the SNR we do not need full tomography of the quantum channel.
In the Supplemental Material we show that the SNR is given by
[…]
×|$\times\absolutevalue{\langle Q\rangle_{\mathcal{T}_{\ket{0}}^{(N)}}+\langle Q\rangle_{\mathcal{T}_{\ket{1}}^{(N)}}-N}$ | (B1)
where
$\Delta K^{2}_{\mathcal{P}_{\ket{j},t}^{(N)}}=$ $\langle Q\rangle_{\mathcal{T}_{\ket{j}}^{(N)}}\Delta K^{2}_{P_{\ket{j},t}}$
$+\left(N-\langle Q\rangle_{\mathcal{T}_{\ket{j}}^{(N)}}\right)\Delta K^{2}_{P_{\ket{\overline{j}},t}}$
$+\left(\langle K\rangle_{P_{\ket{j},t}}-\langle K\rangle_{P_{\ket{\overline{j}},t}}\right)^{2}\Delta Q^{2}_{\mathcal{T}_{\ket{j}}^{(N)}}$
and $\overline{j}=j+1~{}{\rm mod}~{}2$.
Equation (I.2) is very general. Given the time dependent measurement statistics for a single qubit $P_{\ket{j},t}(k)$ and the CNOT noise as described by the probability distributions $\mathcal{T}_{\ket{j}}^{(N)}(q)$, it is possible to calculate the SNR for an $N$ qubit scheme at any time $t$. For example, experimentally we could collect the time dependent measurement statistics of a qubit in the state |$\ket{j}$ to find $P_{\ket{j},t}(k)$ and perform a - relatively limited - version of tomography on the CNOT gates to find the relevant noise distributions $T_{\ket{j}}^{(N)}(q)$. We could then use (I.2) to generate similar plots to Fig. 2 which can be analysed to optimise over measurement duration, number of qubits and SNR.
The full derivation and further discussion of (I.2) can be found in the Supplemental Material.